# Challenges and proposed solutions in modeling multimodal data: A systematic review

### v3

@@ 2 Results
Classification (AUC / Acc) 33, 36, 19, 55 76, 56, 15, 16, 74, 22, 41, 65 77, 64, 58, 23, 44, 81 72, 46, 30, 60, xu_explainable_2021, 14 not reported 33, 22; Acc=97% 19; Acc= 94% 55; Acc= 88% 76; Acc= 0.80 56; Acc=99.1% 15; Acc=99.46% 16; Acc= 97% 74; AUC= 0.86 65; Acc= 91% 77; C-index= 0.86–0.94
64; Acc= 93% 58; Acc= 94% 23; Acc= 75% 44; AUC= 0.90 81; AUC= 0.95 72; AUC= 0.91 xu_explainable_2021; AUC= 0.94 41; Acc= 0.91 46; AUC= 0.9 30
Prediction 38, 39, 57, 63, 59, 69, 47, 49 $wR\approx 0.72\text{--}0.80$ 49; AUROC= 0.92 59;
Not reported 57; AUROC= 0.90 63;
AUC= 0.77 38; ROC $\approx$ 0.90 39; AUC= 0.79 69;
MAPE= 0.165 47
Survival prediction (C-index) 42, 50 AUC= 0.75 42; Acc= 88% 50
Joint (Class. + Regr.) 79 Acc= 0.912 79
Generation / Reconstruction 51 Not reported
@@
Survival prediction (C-index) 68, 20, C-index= 0.63 68; C-index= 0.60–0.75 20
Other (Federated learning) 18 Not reported
@@
Classification (AUC / Acc) 37, 32, Acc= 96.9% 37, AUC= 0.92 32
Biomarker discovery / diagnostics 35 Not reported
@@
Classification (Acc) 58 Acc= 93.4% 58
Representation / Feature learning 54, 31 Acc= 95% 54; Acc= 59% 31
Joint (Class. + Regr.) 29 AUPRC= 0.56 29
Regression (MAE) 66 RMSE reduction=55% 66
@@
Classification (Acc) 67 Acc= 75%
Representation learning (Acc) 78 Acc= 95% (ETH-80);
Prediction (Bias correction) 57 not reported in standard biomedical metrics
Authentication (Error rates) 75 reduced EER by up to 32% and HTER by 28%
@@
Prediction (AUROC) 34, 53, 63 AUROC = 0.84–0.96 34; AUC= 0.94;
F1= 0.88 53, AUROC= 0.85–0.90 63
2.1 Challenges and solutions
2.1.1 Missing of entire modality
@@
In multimodal data settings, there are typically two types of missing data: the first involves missing values within a modality, where individual modalities contain incomplete information, and the second involves the complete absence of an entire modality for certain samples. This means that for some data points, one type of information might be entirely missing, adding an extra layer of complexity. Addressing this challenge requires specialized imputation approaches distinct from those used for handling missing values within individual modalities. Missing modalities are a frequent issue in medical settings, as various examinations are typically conducted for different patients. Moreover, patients may lack access to certain modalities due to factors such as high costs, safety concerns, sensor malfunction, data corruption, or patient dropout 82.
@@
While this systematic review focuses on missing modalities in multimodal data, it’s important to acknowledge the broader literature on missing data, which offers foundational approaches for handling such challenges. Traditional methods for handling missing data involve simple imputation methods 83, which replace missing values with mean or median from the known data, and model-based imputation 84 which uses multivariate statistical methods and leverages feature dependencies to impute missing values. Examples of such model-based methods include k-nearest neighbors 85, MissForest (which uses Random Forests) 86, and matrix completion methods 87. Additionally, multiple imputation methods 88 create several different plausible datasets by filling in missing values multiple times and combining the results to account for the uncertainty of the imputed values. However, in the context of multimodal datasets, these approaches face limitations. Multimodal datasets contain different types of information from various sources for the same subject, and these approaches do not effectively consider the correlations between these modalities (essential information missed in one modality can be maintained in another modality). For instance, simple imputation methods might oversimplify complex relationships present between modalities, model-based imputation relies on assumptions of linear relationships among modalities, and deep learning methods might not explicitly leverage cross-modality correlations. Since each modality contributes unique insights that need to be considered collectively for accurate imputation, it is necessary to develop techniques that can effectively capture and leverage these interdependencies between different modalities.
@@
Common deep learning-based approaches, such as Variational Autoencoder (VAE)-based and Generative Adversarial Network (GAN)-based methods, provide innovative solutions for missing data imputation. However, they face challenges in capturing multimodal correlations. For instance, the VAE model proposed by McCoy et al. 89 can handle missing data but does not account for cross-modality relationships. To address this, Wu and Goodman 90 extended the VAE framework to the Multimodal Variational Autoencoder (MVAE), which utilizes a product of expert inference networks to better integrate unique modality contributions 91. However, MVAE struggles to handle randomly missing modalities effectively due to its reliance on fixed input-output structures, limiting its use in clinical applications.
@@
For taking the correlation between modalities into account, Saad et al. 43 proposed using a multitasking learning approach. This approach involves dividing incomplete multimodal datasets into subsets, each containing a different combination of modalities. Individual classifiers are then trained independently on these subsets, and their outputs are combined to generate the final classifier. However, this method assumes linear relationships between modalities, which is often unrealistic. Another approach, proposed by Vivar et al. 72, involves matrix completion techniques by concatenating multimodal features into a single matrix. Yet, this approach fails when entire modalities are missing, as missing entries form blocks rather than random distributions, violating assumptions of matrix completion methods. To address these limitations, extensions of VAE and GAN frameworks have been proposed. For example, Xu et al. 48 introduced the Dynamic Multimodal Variational Autoencoder (DMVAE), which incorporates modality indicators for missing modalities and leverages joint representations to capture correlations across modalities. This enables DMVAE to impute missing features more effectively than MVAE. However, DMVAE is not well-suited for image data. In this context, GAN-based approaches provide alternative solutions. Cai et al. 71 adapted GANs for multimodal settings, using an encoder-decoder architecture to generate missing modalities based on existing ones. To extend this capability beyond imaging, Ziegler et al. 51 proposed the Multimodal Conditional GAN (MMCGAN). This model combines conditional GANs for tabular data with models for conditional 3D image synthesis, preserving cross-modal correlations and generating realistic synthetic data across diverse modalities. In addition to these generative methods, Hou et al. 20 introduced a Hybrid Graph Convolutional Network that incorporates an online masked autoencoder to leverage cross-modal dependencies via transformers. However Wang et al. 67 highlighted a drawback of imputation-based methods, noting that they can introduce noise, especially when only a limited number of complete modality samples are available, which can negatively affect downstream tasks. To address this, they proposed a novel multimodal learning framework that leverages all available samples, including those with incomplete modalities, without relying on imputation. Their approach is based on knowledge distillation 92. In this framework, modality-specific models are first trained independently on all available data, even if some modalities are missing. These models act as “teachers” to train a multimodal “student” model. The student model learns by combining the soft labels provided by the teacher models with the true one-hot labels. Because each teacher model is trained exclusively on a single modality, it benefits from a larger effective sample size, making it a robust expert for its respective modality. The student model integrates these modality-specific insights, enabling it to combine information across all modalities effectively. This approach stands out because it does not discard incomplete modality samples or directly impute missing data. Instead, it ensures that even incomplete samples contribute to the expertise of the teacher models, maximizing data utilization 67. However, Du et al. 93 noted that not all knowledge sources are equally reliable, and conflicting information can arise during distillation. Simply weighing all knowledge sources equally can reduce the effectiveness of the student model. To address this limitation, Xing et al. 32 proposed the Gradient-guided Knowledge Modulation (GKM) scheme. GKM dynamically adjusts the influence of each knowledge source based on its consistency in the gradient space. By prioritizing reliable information and filtering out misleading knowledge, GKM enhances the distillation process, ensuring the student model absorbs useful insights while minimizing conflicts. This selective adjustment improves the robustness and performance of the student network, offering more dependable guidance from diverse sources.
2.1.2 Small dataset
@@
One of the major challenges in multimodal data fusion is dealing with small data. In total, 17 publications in this review tackled this challenge, reflecting its critical importance in multimodal modeling Limited, insufficient, or sparse data in multimodal fusion complicates the integration of multiple modalities, making it challenging to capture the true underlying patterns and relationships. Researchers have proposed various innovative approaches to address the small data challenge, ranging from data augmentation to transfer learning. Data augmentation techniques 15 stand out, artificially expanding dataset size and diversity, thereby enhancing the training of robust fusion models. Another strategy involves multimodal pre-training, where models learn generic representations from large unimodal datasets, applicable to smaller multimodal datasets. Multi-task learning 94, 64 emerges as a powerful technique, enabling the model to jointly train on related tasks, leveraging shared information across modalities to improve performance, even with limited data. Ensemble learning, combining multiple models or techniques, also proves effective in addressing the small data challenge. On the other hand, Cai et al. 36 proposed a novel multimodal transformer for image and metadata fusion, incorporating pre-trained Vision Transformer (ViT) models.
@@
In the context of medical imaging, Li et al. 22 proposes transfer learning methods, involving fine-tuning CNN models pre-trained on natural image datasets to overcome the scarcity of medical image data. Similarly, Shickel et al. 59, Lim et al. 41, Jabeen et al. 15 and Yoo et al. 50 overcome small sample size challenge through transfer learning. This involves initializing a smaller model by leveraging pre-existing models trained on larger datasets. These initialized models are subsequently fine-tuned using the smaller, domain-specific dataset, thereby optimizing performance and generalization for medical image analysis tasks
@@
However, Lei et al. 74 argued that the “small-n-large-p” problem, characterized by a large feature dimension and small sample size, presents persistent challenges in identifying clinical subjects correctly via robust modeling. Both dimension reduction and feature selection offer promising solutions, addressing over-fitting issues and reducing computational time. Saad et al. 43, taking a different approach to tackle this challenge, they utilize traditional canonical correlation analysis for data fusion highlighting its efficacy in leveraging hidden linear correlation among features that are both small size and heterogeneous. Which can overcome the challenge of small datasets, particularly for 2D images and 1D genomic data.
@@
Kustowski et al. 57 argued against a purely data-driven approach with limited data and proposed a hybrid methodology combining knowledge-driven and data-driven approaches. The process begins with a simulation-trained neural network surrogate model, which captures crucial correlations among different data modalities and between simulation inputs and outputs. Subsequently, this knowledge is transferred to retrain the model, aligning it with real-world data for improved predictive performance. In another approach, Guan et al. 56 addressed the challenge of limited data by using knowledge distillation. This technique transfers insights from a more complex teacher network to a smaller student network 95. In a similar way, the knowledge obtained from multimodal data by a teacher model can be transferred to a student model. Faced with a scarcity of Magnetic Resonance Imaging (MRI) data, Guan et al. 56 propose a multi-modal multi-instance distillation strategy. This approach aims to transfer insights from a multi-modal teacher network to a single-modal student network. The teacher network, trained with both MRI and clinical data, captured comprehensive disease-related patterns. This knowledge was distilled into the student network, which was trained only on MRI data, ensuring that the richer, multi-modal insights were preserved and utilized effectively.
2.1.3 Interpretability
@@
Interpretability in multimodal modeling refers to understanding how a model integrates and assigns significance to information from various sources, making its predictions explainable. This is particularly important in medical fields, where decisions made based on model outputs can directly impact patient health. In this review, 14 publications address the challenge of interpretability, highlighting its essential role in supporting model reliability and understanding in healthcare applications. Various techniques have been developed to improve the interpretability of deep learning models in multimodal environments. One such method is the use of attention mechanisms, which allow the model to focus on specific aspects of the data during prediction. For instance, these mechanisms can highlight important regions in a medical image or key elements in genomic sequences, thereby improving transparency in the prediction process. Another approach involves gradient-based attribution algorithms, which identify and quantify the contribution of each modality to the final prediction. These methods are particularly useful in understanding the biological relevance of the model’s outputs. For instance, to increase the interpretability in a classification task and fusion of metadata and medical imaging data 22 introduced the “gradient-weighted class activation mapping” method, which generates visual explanations by highlighting the regions of the input medical images that are important for the model’s prediction.
@@
For the fusion of multi-omic data and clinical data within the context of prognosis prediction in survival analysis, Jiang et al. 96 and Hao et al. 68 addressed the challenge of the black-box nature of deep neural networks in two different ways. Jiang et al. 96 incorporated Deep SHapley Additive exPlanations (DeepSHAP), a technique that combines deep learning with Shapley values, to explain individual predictions made by the model. The AUTOSurv model, which predicts cancer prognosis, uses DeepSHAP to analyze the SHAP values and quantify how each input feature, such as gene expression or miRNA data, contributes to the final output. This approach enhances the model’s interpretability by linking key pathways with relevant genes, which not only improves the model’s transparency but also highlights potential biomarkers for cancer prognosis.
@@
Similarly, Hao et al. 68 developed CoxPSANet, a deep learning model for cancer survival prediction, which integrates gene expression data at the individual gene level and biological pathways. By establishing sparse connections between the gene and pathway layers, CoxPSANet enhances interpretability by focusing on key biological components and pathways that are relevant to cancer prognosis. For the fusion of imaging and non-imaging data, Xin et al. 60 proposed the “Interpretable Deep Multimodal Fusion” framework, which utilizes Canonical Correlation Analysis to interpret the complex, non-linear associations between modalities. They introduced a cross-modal association (CA) score to assess the impact of each feature, further enhancing the interpretability of the multimodal fusion process.
2.1.4 Imbalance in dimensionality
@@
In our review, 7 studies addressed this challenge, using various techniques. For example, Karaman et al. 39 and Partin et al. 25 employed weighted loss functions to assign greater importance to underrepresented modalities or classes, counteracting the overemphasis on high-dimensional data. Similarly, Mao et al. 58 and Li et al. 22 utilized focal loss functions to concentrate learning on harder-to-classify samples, which helped reduce bias toward dominant modalities and enhanced model robustness. Other approaches included non-linear dimensionality reduction, as demonstrated by Madabhushi et al. 80, which aimed to preserve the intrinsic structure of the data while minimizing dimensionality-related imbalances. In a different vein, Xu et al. 49 leveraged recurrent neural networks to model temporal dependencies, helping to better align and integrate multimodal inputs over time.
@@
Although dimensionality imbalance focuses on feature-level disparities, related issues such as class imbalance, particularly common in high-dimensional spaces, can further complicate multimodal classification tasks. Tran and Liatsis 75 addressed this by introducing the RABOC algorithm, a hybrid method combining one-class classification and Real AdaBoost. One-class classification is inherently suited to extreme class imbalance, as it defines decision boundaries using only the target class, while Real AdaBoost boosts performance by focusing on misclassified samples through weighted re-training. Together, these methods effectively handle severe class imbalance in high-dimensional multimodal biometric authentication tasks without overfitting. Similarly, Li et al. 22 tackled class imbalance in multimodal classification through the Deep Multimodal Generative Adversarial Network (DMGAN). This approach generates synthetic features for minority classes via adversarial training, thereby rebalancing the dataset. DMGAN outperforms traditional oversampling techniques like SMOTE (Synthetic Minority Over-sampling Technique) 97, XGBoost (eXtreme Gradient Boosting) 98, Unimodal GAN frameworks 99, and Random Forest 100, demonstrating the effectiveness of generative approaches in dealing with imbalanced multimodal data.
2.1.5 Optimal fusion technique
@@
To make the most out of multimodal data, an optimal fusion strategy is required. The optimal fusion strategy in modeling multimodal data depends on various factors such as the specific task requirements, the characteristics of the modalities involved (e.g., the number of modalities, their relationships, dimensionality, and sampling rates), the available computational resources, trade-offs between model complexity, interpretability, performance, and the ease of implementation 101. Additionally, the selection of an optimal fusion strategy should also consider the quality and reliability of each modality’s data, as well as the level of correlation between the modalities 30. By considering all these factors, the most effective fusion strategy can be determined for specific multimodal data modeling tasks. Overall, an optimal fusion strategy in modeling multimodal data involves integrating and combining multiple modalities in a way that maximizes the benefits of each modality while minimizing the biases and limitations inherent in individual modalities 5.

### v4

@@ Results
Classification (AUC / Acc) Aghili et al. [2022], Cai et al. [2022], Dai et al. [2023], Dong et al. [2021], Eshaghi et al. [2015], Guan et al. [2021], Jabeen et al. [2022], Khan et al. [2020], Lei et al. [2016], Li et al. [2023], Lim et al. [2022], Liu et al. [2020, 2014, 2020], Mao et al. [2021], Martí-Juan et al. [2023], Sheu et al. [2022], Tiwari et al. [2011], Vivar et al. [2018], Wang et al. [2022, 2023], Xin et al. [2021], Xu et al. [2021], Zhang et al. [2023] not reported Aghili et al. [2022], Li et al. [2023]; Acc=97% Dai et al. [2023]; Acc= 94% Dong et al. [2021]; Acc= 88% Eshaghi et al. [2015]; Acc= 0.80 Guan et al. [2021]; Acc=99.1% Jabeen et al. [2022]; Acc=99.46% Khan et al. [2020]; Acc= 97% Lei et al. [2016]; AUC= 0.86 Liu et al. [2020]; Acc= 91% Liu et al. [2014]; C-index= 0.86–0.94
Liu et al. [2020]; Acc= 93% Mao et al. [2021]; Acc= 94% Martí-Juan et al. [2023]; Acc= 75% Sheu et al. [2022]; AUC= 0.90 Tiwari et al. [2011]; AUC= 0.95 Vivar et al. [2018]; AUC= 0.91 Xu et al. [2021]; AUC= 0.94 Lim et al. [2022]; Acc= 0.91 Wang et al. [2022]; AUC= 0.9 Wang et al. [2023]
Prediction Hatami et al. [2022], Karaman et al. [2022], Kustowski et al. [2021], Li et al. [2020], Shickel et al. [2021], Tiulpin et al. [2019], Wang et al. [2022], Xu et al. [2022] $wR\approx 0.72\text{--}0.80$ Xu et al. [2022]; AUROC= 0.92 Shickel et al. [2021];
Not reported Kustowski et al. [2021]; AUROC= 0.90 Li et al. [2020];
AUC= 0.77 Hatami et al. [2022]; ROC $\approx$ 0.90 Karaman et al. [2022]; AUC= 0.79 Tiulpin et al. [2019];
MAPE= 0.165 Wang et al. [2022]
Survival prediction (C-index) Rabinovici-Cohen et al. [2022], Yoo et al. [2022] AUC= 0.75 Rabinovici-Cohen et al. [2022]; Acc= 88% Yoo et al. [2022]
Joint (Class. + Regr.) Thung et al. [2013] Acc= 0.912 Thung et al. [2013]
Generation / Reconstruction Ziegler et al. [2022] Not reported
@@
Survival prediction (C-index) Hao et al. [2019], Hou et al. [2023], C-index= 0.63 Hao et al. [2019]; C-index= 0.60–0.75 Hou et al. [2023]
Other (Federated learning) Cremonesi et al. [2023] Not reported
@@
Classification (AUC / Acc) Golovanevsky et al. [2022], Xing et al. [2023], Acc= 96.9% Golovanevsky et al. [2022], AUC= 0.92 Xing et al. [2023]
Biomarker discovery / diagnostics Boehm et al. [2022] Not reported
@@
Classification (Acc) Mao et al. [2021] Acc= 93.4% Mao et al. [2021]
Representation / Feature learning Bahador et al. [2021], Xia et al. [2023] Acc= 95% Bahador et al. [2021]; Acc= 59% Xia et al. [2023]
Joint (Class. + Regr.) Sükei et al. [2023] AUPRC= 0.56 Sükei et al. [2023]
Regression (MAE) von Lühmann et al. [2020] RMSE reduction=55% von Lühmann et al. [2020]
@@
Classification (Acc) Wang et al. [2020] Acc= 75%
Representation learning (Acc) Zhang et al. [2014] Acc= 95% (ETH-80);
Prediction (Bias correction) Kustowski et al. [2021] not reported in standard biomedical metrics
Authentication (Error rates) Tran and Liatsis [2016] reduced EER by up to 32% and HTER by 28%
@@
Prediction (AUROC) Ahuja et al. [2022], An et al. [2021], Li et al. [2020] AUROC = 0.84–0.96 Ahuja et al. [2022]; AUC= 0.94;
F1= 0.88 An et al. [2021], AUROC= 0.85–0.90 Li et al. [2020]
Challenges and solutions
Missing of entire modality
@@
In multimodal data settings, there are typically two types of missing data: the first involves missing values within a modality, where individual modalities contain incomplete information, and the second involves the complete absence of an entire modality for certain samples. This means that for some data points, one type of information might be entirely missing, adding an extra layer of complexity. Addressing this challenge requires specialized imputation approaches distinct from those used for handling missing values within individual modalities. Missing modalities are a frequent issue in medical settings, as various examinations are typically conducted for different patients. Moreover, patients may lack access to certain modalities due to factors such as high costs, safety concerns, sensor malfunction, data corruption, or patient dropout Zhang et al. [2022].
@@
While this systematic review focuses on missing modalities in multimodal data, it’s important to acknowledge the broader literature on missing data, which offers foundational approaches for handling such challenges. Traditional methods for handling missing data involve simple imputation methods Little and Rubin [2019], which replace missing values with mean or median from the known data, and model-based imputation Ngueilbaye et al. [2021] which uses multivariate statistical methods and leverages feature dependencies to impute missing values. Examples of such model-based methods include k-nearest neighbors Liao et al. [2014], MissForest (which uses Random Forests) Tang and Ishwaran [2017], and matrix completion methods Fan et al. [2020]. Additionally, multiple imputation methods Rubin [2018] create several different plausible datasets by filling in missing values multiple times and combining the results to account for the uncertainty of the imputed values. However, in the context of multimodal datasets, these approaches face limitations. Multimodal datasets contain different types of information from various sources for the same subject, and these approaches do not effectively consider the correlations between these modalities (essential information missed in one modality can be maintained in another modality). For instance, simple imputation methods might oversimplify complex relationships present between modalities, model-based imputation relies on assumptions of linear relationships among modalities, and deep learning methods might not explicitly leverage cross-modality correlations. Since each modality contributes unique insights that need to be considered collectively for accurate imputation, it is necessary to develop techniques that can effectively capture and leverage these interdependencies between different modalities.
@@
Common deep learning-based approaches, such as Variational Autoencoder (VAE)-based and Generative Adversarial Network (GAN)-based methods, provide innovative solutions for missing data imputation. However, they face challenges in capturing multimodal correlations. For instance, the VAE model proposed by McCoy et al. [2018] can handle missing data but does not account for cross-modality relationships. To address this, Wu and Goodman [2018] extended the VAE framework to the Multimodal Variational Autoencoder (MVAE), which utilizes a product of expert inference networks to better integrate unique modality contributions Hinton [2002]. However, MVAE struggles to handle randomly missing modalities effectively due to its reliance on fixed input-output structures, limiting its use in clinical applications.
@@
For taking the correlation between modalities into account, Saad et al. [2022] proposed using a multitasking learning approach. This approach involves dividing incomplete multimodal datasets into subsets, each containing a different combination of modalities. Individual classifiers are then trained independently on these subsets, and their outputs are combined to generate the final classifier. However, this method assumes linear relationships between modalities, which is often unrealistic. Another approach, proposed by Vivar et al. [2018], involves matrix completion techniques by concatenating multimodal features into a single matrix. Yet, this approach fails when entire modalities are missing, as missing entries form blocks rather than random distributions, violating assumptions of matrix completion methods. To address these limitations, extensions of VAE and GAN frameworks have been proposed. For example, Xu et al. [2022] introduced the Dynamic Multimodal Variational Autoencoder (DMVAE), which incorporates modality indicators for missing modalities and leverages joint representations to capture correlations across modalities. This enables DMVAE to impute missing features more effectively than MVAE. However, DMVAE is not well-suited for image data. In this context, GAN-based approaches provide alternative solutions. Cai et al. [2018] adapted GANs for multimodal settings, using an encoder-decoder architecture to generate missing modalities based on existing ones. To extend this capability beyond imaging, Ziegler et al. [2022] proposed the Multimodal Conditional GAN (MMCGAN). This model combines conditional GANs for tabular data with models for conditional 3D image synthesis, preserving cross-modal correlations and generating realistic synthetic data across diverse modalities. In addition to these generative methods, Hou et al. [2023] introduced a Hybrid Graph Convolutional Network that incorporates an online masked autoencoder to leverage cross-modal dependencies via transformers. However Wang et al. [2020] highlighted a drawback of imputation-based methods, noting that they can introduce noise, especially when only a limited number of complete modality samples are available, which can negatively affect downstream tasks. To address this, they proposed a novel multimodal learning framework that leverages all available samples, including those with incomplete modalities, without relying on imputation. Their approach is based on knowledge distillation Hinton et al. [2015]. In this framework, modality-specific models are first trained independently on all available data, even if some modalities are missing. These models act as “teachers” to train a multimodal “student” model. The student model learns by combining the soft labels provided by the teacher models with the true one-hot labels. Because each teacher model is trained exclusively on a single modality, it benefits from a larger effective sample size, making it a robust expert for its respective modality. The student model integrates these modality-specific insights, enabling it to combine information across all modalities effectively. This approach stands out because it does not discard incomplete modality samples or directly impute missing data. Instead, it ensures that even incomplete samples contribute to the expertise of the teacher models, maximizing data utilization Wang et al. [2020]. However, Du et al. [2020] noted that not all knowledge sources are equally reliable, and conflicting information can arise during distillation. Simply weighing all knowledge sources equally can reduce the effectiveness of the student model. To address this limitation, Xing et al. [2023] proposed the Gradient-guided Knowledge Modulation (GKM) scheme. GKM dynamically adjusts the influence of each knowledge source based on its consistency in the gradient space. By prioritizing reliable information and filtering out misleading knowledge, GKM enhances the distillation process, ensuring the student model absorbs useful insights while minimizing conflicts. This selective adjustment improves the robustness and performance of the student network, offering more dependable guidance from diverse sources.
Small dataset
@@
One of the major challenges in multimodal data fusion is dealing with small data. In total, 17 publications in this review tackled this challenge, reflecting its critical importance in multimodal modeling Limited, insufficient, or sparse data in multimodal fusion complicates the integration of multiple modalities, making it challenging to capture the true underlying patterns and relationships. Researchers have proposed various innovative approaches to address the small data challenge, ranging from data augmentation to transfer learning. Data augmentation techniques Jabeen et al. [2022] stand out, artificially expanding dataset size and diversity, thereby enhancing the training of robust fusion models. Another strategy involves multimodal pre-training, where models learn generic representations from large unimodal datasets, applicable to smaller multimodal datasets. Multi-task learning Saad et al. [2022], Liu et al. [2020] emerges as a powerful technique, enabling the model to jointly train on related tasks, leveraging shared information across modalities to improve performance, even with limited data. Ensemble learning, combining multiple models or techniques, also proves effective in addressing the small data challenge. On the other hand, Cai et al. [2022] proposed a novel multimodal transformer for image and metadata fusion, incorporating pre-trained Vision Transformer (ViT) models.
@@
In the context of medical imaging, Li et al. [2023] proposes transfer learning methods, involving fine-tuning CNN models pre-trained on natural image datasets to overcome the scarcity of medical image data. Similarly, Shickel et al. [2021], Lim et al. [2022], Jabeen et al. [2022] and Yoo et al. [2022] overcome small sample size challenge through transfer learning. This involves initializing a smaller model by leveraging pre-existing models trained on larger datasets. These initialized models are subsequently fine-tuned using the smaller, domain-specific dataset, thereby optimizing performance and generalization for medical image analysis tasks
@@
However, Lei et al. [2016] argued that the “small-n-large-p” problem, characterized by a large feature dimension and small sample size, presents persistent challenges in identifying clinical subjects correctly via robust modeling. Both dimension reduction and feature selection offer promising solutions, addressing over-fitting issues and reducing computational time. Saad et al. [2022], taking a different approach to tackle this challenge, they utilize traditional canonical correlation analysis for data fusion highlighting its efficacy in leveraging hidden linear correlation among features that are both small size and heterogeneous. Which can overcome the challenge of small datasets, particularly for 2D images and 1D genomic data.
@@
Kustowski et al. [2021] argued against a purely data-driven approach with limited data and proposed a hybrid methodology combining knowledge-driven and data-driven approaches. The process begins with a simulation-trained neural network surrogate model, which captures crucial correlations among different data modalities and between simulation inputs and outputs. Subsequently, this knowledge is transferred to retrain the model, aligning it with real-world data for improved predictive performance. In another approach, Guan et al. [2021] addressed the challenge of limited data by using knowledge distillation. This technique transfers insights from a more complex teacher network to a smaller student network Gou et al. [2021]. In a similar way, the knowledge obtained from multimodal data by a teacher model can be transferred to a student model. Faced with a scarcity of Magnetic Resonance Imaging (MRI) data, Guan et al. [2021] propose a multi-modal multi-instance distillation strategy. This approach aims to transfer insights from a multi-modal teacher network to a single-modal student network. The teacher network, trained with both MRI and clinical data, captured comprehensive disease-related patterns. This knowledge was distilled into the student network, which was trained only on MRI data, ensuring that the richer, multi-modal insights were preserved and utilized effectively.
Interpretability
@@
Interpretability in multimodal modeling refers to understanding how a model integrates and assigns significance to information from various sources, making its predictions explainable. This is particularly important in medical fields, where decisions made based on model outputs can directly impact patient health. In this review, 14 publications address the challenge of interpretability, highlighting its essential role in supporting model reliability and understanding in healthcare applications. Various techniques have been developed to improve the interpretability of deep learning models in multimodal environments. One such method is the use of attention mechanisms, which allow the model to focus on specific aspects of the data during prediction. For instance, these mechanisms can highlight important regions in a medical image or key elements in genomic sequences, thereby improving transparency in the prediction process. Another approach involves gradient-based attribution algorithms, which identify and quantify the contribution of each modality to the final prediction. These methods are particularly useful in understanding the biological relevance of the model’s outputs. For instance, to increase the interpretability in a classification task and fusion of metadata and medical imaging data Li et al. [2023] introduced the “gradient-weighted class activation mapping” method, which generates visual explanations by highlighting the regions of the input medical images that are important for the model’s prediction.
@@
For the fusion of multi-omic data and clinical data within the context of prognosis prediction in survival analysis, Jiang et al. [2024] and Hao et al. [2019] addressed the challenge of the black-box nature of deep neural networks in two different ways. Jiang et al. [2024] incorporated Deep SHapley Additive exPlanations (DeepSHAP), a technique that combines deep learning with Shapley values, to explain individual predictions made by the model. The AUTOSurv model, which predicts cancer prognosis, uses DeepSHAP to analyze the SHAP values and quantify how each input feature, such as gene expression or miRNA data, contributes to the final output. This approach enhances the model’s interpretability by linking key pathways with relevant genes, which not only improves the model’s transparency but also highlights potential biomarkers for cancer prognosis.
@@
Similarly, Hao et al. [2019] developed CoxPSANet, a deep learning model for cancer survival prediction, which integrates gene expression data at the individual gene level and biological pathways. By establishing sparse connections between the gene and pathway layers, CoxPSANet enhances interpretability by focusing on key biological components and pathways that are relevant to cancer prognosis. For the fusion of imaging and non-imaging data, Xin et al. [2021] proposed the “Interpretable Deep Multimodal Fusion” framework, which utilizes Canonical Correlation Analysis to interpret the complex, non-linear associations between modalities. They introduced a cross-modal association (CA) score to assess the impact of each feature, further enhancing the interpretability of the multimodal fusion process.
Imbalance in dimensionality
@@
In our review, 7 studies addressed this challenge, using various techniques. For example, Karaman et al. [2022] and Partin et al. [2023] employed weighted loss functions to assign greater importance to underrepresented modalities or classes, counteracting the overemphasis on high-dimensional data. Similarly, Mao et al. [2021] and Li et al. [2023] utilized focal loss functions to concentrate learning on harder-to-classify samples, which helped reduce bias toward dominant modalities and enhanced model robustness. Other approaches included non-linear dimensionality reduction, as demonstrated by Madabhushi et al. [2011], which aimed to preserve the intrinsic structure of the data while minimizing dimensionality-related imbalances. In a different vein, Xu et al. [2022] leveraged recurrent neural networks to model temporal dependencies, helping to better align and integrate multimodal inputs over time.
@@
Although dimensionality imbalance focuses on feature-level disparities, related issues such as class imbalance, particularly common in high-dimensional spaces, can further complicate multimodal classification tasks. Tran and Liatsis [2016] addressed this by introducing the RABOC algorithm, a hybrid method combining one-class classification and Real AdaBoost. One-class classification is inherently suited to extreme class imbalance, as it defines decision boundaries using only the target class, while Real AdaBoost boosts performance by focusing on misclassified samples through weighted re-training. Together, these methods effectively handle severe class imbalance in high-dimensional multimodal biometric authentication tasks without overfitting. Similarly, Li et al. [2023] tackled class imbalance in multimodal classification through the Deep Multimodal Generative Adversarial Network (DMGAN). This approach generates synthetic features for minority classes via adversarial training, thereby rebalancing the dataset. DMGAN outperforms traditional oversampling techniques like SMOTE (Synthetic Minority Over-sampling Technique) Alex and Nayahi [2023], XGBoost (eXtreme Gradient Boosting) Chen and Guestrin [2016], Unimodal GAN frameworks Sauber-Cole and Khoshgoftaar [2022], and Random Forest More and Rana [2017], demonstrating the effectiveness of generative approaches in dealing with imbalanced multimodal data.
Optimal fusion technique
@@
To make the most out of multimodal data, an optimal fusion strategy is required. The optimal fusion strategy in modeling multimodal data depends on various factors such as the specific task requirements, the characteristics of the modalities involved (e.g., the number of modalities, their relationships, dimensionality, and sampling rates), the available computational resources, trade-offs between model complexity, interpretability, performance, and the ease of implementation Ramachandram and Taylor [2017]. Additionally, the selection of an optimal fusion strategy should also consider the quality and reliability of each modality’s data, as well as the level of correlation between the modalities Wang et al. [2023]. By considering all these factors, the most effective fusion strategy can be determined for specific multimodal data modeling tasks. Overall, an optimal fusion strategy in modeling multimodal data involves integrating and combining multiple modalities in a way that maximizes the benefits of each modality while minimizing the biases and limitations inherent in individual modalities Pawłowski et al. [2023].